OpenStudy (anonymous):

I FAN AND GIVE MEDALS. Simplify a^9/a^4

OpenStudy (anonymous):

all you have to do is subtract 9 - 4 = 5 a^5

OpenStudy (anonymous):

Oh yeah... okay I get that, can you help me with a few more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

alright uhhhhh im pretty sure all you have to do is subtract the like terms like subtract the r^3 from the r^3, s^2 from s^4, and one t because thats all they have in common so itd be s^2t^3? but im not too sure you might wanna look it up in your textbook or s/t

OpenStudy (anonymous):

Okay yeah thats what I'm thinking too

OpenStudy (anonymous):

do you have any more?

OpenStudy (anonymous):

nvm prematurely posted it ok for this one you have to simplify it 4/16 simplified is 1/4 and then do what you did in the last problem (subtracting the bottom from the top) 1/4x^6

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

nope! im real sorry hon try mathway.com maybe?

OpenStudy (anonymous):

No it's okay :D Identify the property demonstrated: 4^3 • 4^7 = 4^3+7

OpenStudy (anonymous):

A. Quotient of Powers Property B. Product of Powers Property C. Power of a Product Property D. Power of a Power Property

OpenStudy (anonymous):

well theres multiplication involved so you can rule out A D sounds stupid id go with C but im not sure

OpenStudy (anonymous):

Kay I have one more? Identify the property demonstrated: (5^3)^2 = 5^3 • 2 A. Quotient of Powers Property B. Product of Powers Property C. Power of a Power Property D. Power of a Product Property

OpenStudy (anonymous):

id maybe say d im just as clueless as you though i know im gonna sound like no fun but try lookin through your lessons again? just like look over what those four properties mean or google em or s/t

OpenStudy (anonymous):

cool thanks :)

OpenStudy (anonymous):

anytime yo

OpenStudy (anonymous):

:*